OpenStudy (anonymous):

sin x /1 - cos x + sin x / 1- cos x = 2 csc x

OpenStudy (zzr0ck3r):

\[\frac{2sin(x)}{1-cos(x)}=\frac{2}{sin(x)}\]

OpenStudy (zzr0ck3r):

with me?

OpenStudy (anonymous):

No how did you get 2 sin x

OpenStudy (anonymous):

divided by 1 -cos x

OpenStudy (anonymous):

You still here ?

OpenStudy (anonymous):

Its proving trig equations if that helps solve it

OpenStudy (zzr0ck3r):

sorry had to run to the store \[\frac{sin(x)}{1-cos(x)}+\frac{sin(x)}{1-cos(x)}\]

OpenStudy (zzr0ck3r):

is this what you have?

OpenStudy (zzr0ck3r):

@Seadog12 ?

OpenStudy (anonymous):

yeah but the denominator of the last one is 1 + cosx

OpenStudy (zzr0ck3r):

ahhh \[\frac{sinx}{1-cos(x)}+\frac{sinx}{1+cosx}\] find common denominator \[\frac{sin(x)(1+cos(x))+sin(x)(1-cos(x))}{(1-cos(x))(1+cos(x))}=\\\frac{sin(x)+sin(x)cos(x)+sin(x)-cos(x)sin(x)}{(1-cos^2(x))}=\frac{2sin(x)}{sin^2(x)}=\frac{2}{sin(x)}=\\2csc(x)\]

OpenStudy (zzr0ck3r):

understand?

OpenStudy (anonymous):

so the cos cancel out on the top

OpenStudy (anonymous):

the cosines cancel out

OpenStudy (zzr0ck3r):

well look at the second to last line sin(x)+sin(x)cos(x)+sin(x)−cos(x)sin(x) rearrange sin(x)cos(x) - sin(x)cos(x) + sin(x) + sin(x) (sin(x)cos(x)-sin(x)cos(x)) + 2sin(x) = ( 0 ) + 2sin(x) = 2sin(x)

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

Thanks zzr0ck3r